OpenStudy (anonymous):

Suppose f(x)=4x-2 and g(x)=2x+1..... Find the value of f(2)/g(4)

OpenStudy (anonymous):

plug 2 in for x in f(x) and 4 for x in g(x) so youd have f(2)=(4x2)-2=6 g(4)=(2x4)+1=9 f(2)/g(4)=6/9=2/3